OpenStudy (anonymous):

A parallelogram has sides measuring 16.6 m and 33.2 m. The height corresponding to the 16.6-m base is 11.9 m. Find the height,to the nearest tenth of a meter, corresponding to the 33.2-m base.

jimthompson5910 (jim_thompson5910):

Hint: you use the formula A = b*h b = base h = height to find the area of the parallelogram then you use this area to find the height you want

OpenStudy (anonymous):

11.9*16.6?

jimthompson5910 (jim_thompson5910):

multiply those out to get ____ for the area

OpenStudy (anonymous):

197.54 is my area now how do i get the height?

jimthompson5910 (jim_thompson5910):

A = b*h ... area formula 197.54 = b*h ... plug in the area you just found 197.54 = 33.2*h ... plug in the base (you want to find the corresponding height for) now solve for h

OpenStudy (anonymous):

5.95 is the answer?

jimthompson5910 (jim_thompson5910):

correct